OpenStudy (anonymous):

Given: Line PB tangent Line PV, Line PU Secants If m Arch VU = 70° and m Arch ST = 30°, then m Angle 2 = a) 20 b) 35 c) 50

OpenStudy (anonymous):

The answer I got wrong was 35

OpenStudy (phi):

average of the two arcs

OpenStudy (anonymous):

So, add them together and divide by two?

OpenStudy (phi):

yes

OpenStudy (anonymous):

Okay sorry it took so long

OpenStudy (anonymous):

I was Driving off a stalker

OpenStudy (anonymous):

I don't think you wanna' know what he said to me either

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

I'm not worried or anything. Just annoyed

OpenStudy (anonymous):

@phi

OpenStudy (phi):

Report them if it happens again

OpenStudy (anonymous):

I will

OpenStudy (anonymous):

Don't worry

OpenStudy (anonymous):

Oh and btw, there is another question that applies to the same picture

OpenStudy (anonymous):

@phi

OpenStudy (phi):

if it asks for angle 1, that is 1/2 the difference of the two arcs i.e. (big - small)/2

OpenStudy (anonymous):

If m Angle 1 = 30° and m Arch ST = 20°, then m Arch VU = a) 10 b) 40 c) 80

OpenStudy (phi):

\[ \frac{VU- ST}{2} = Angle 1\]

OpenStudy (phi):

fill in what you know, then use algebra to "solve for VU" can you do that ?

OpenStudy (anonymous):

Hmmm. Let me think for a sec

OpenStudy (phi):

in the formula, replace ST with 20 and angle 1 with 30

OpenStudy (phi):

\[ \frac{VU- 20}{2} = 30 \] I would start by multiplying both sides by 2

OpenStudy (anonymous):

Yeah, I'm writing it down

OpenStudy (anonymous):

Oh and thanks. That helps

OpenStudy (anonymous):

I was thinking about doing that, but then I started to second guess myself

OpenStudy (anonymous):

So VU-40= 60?

OpenStudy (anonymous):

VU=100?

OpenStudy (phi):

you do this \[2\cdot \frac{(VU- 20)}{2} = 30 \cdot 2\] you can think of the left side as multiplying fraction times fraction (top times top and bottom times bottom) \[ \frac{2}{1} \cdot \frac{(VU- 20)}{2} = 30 \cdot 2\] \[\frac{2(VU- 20)}{2} = 30 \cdot 2\] or \[ \frac{2}{2} \frac{(VU- 20)}{1} = 30 \cdot 2\]

OpenStudy (phi):

in other words you should get VU -20 = 60

OpenStudy (anonymous):

Okay. So you don't multiply 20 by 2?

OpenStudy (anonymous):

I'm guessing that's where I messed up

OpenStudy (phi):

remember to how to add or subtract fractions: \[ \frac{VU}{2} - \frac{20}{2} = \frac{VU-20}{2} \] you can do it the other way round: \[ \frac{VU-20}{2}= \frac{VU}{2} - \frac{20}{2} \] and simplify 20/2 to 10 \[ \frac{VU}{2} - 10 = 30 \] now add 10 to both sides \[ \frac{VU}{2} - 10 +10 = 30+10 \\ \frac{VU}{2} = 40 \]

OpenStudy (anonymous):

So VU= 80

OpenStudy (phi):

yes.

OpenStudy (anonymous):

Alright I understand it now

OpenStudy (phi):

I guess you have to get the idea that \[ \frac{complicated\ stuff}{2} \cdot 2 = complicated\ stuff \]

OpenStudy (anonymous):

Hehehe. Lol!

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

I'm gonna' be quoting that for awhile

OpenStudy (anonymous):

Thanks!

OpenStudy (phi):

yw